The Neal Company wants to estimate next years return on equity (ROE) under different financial leverage ratios. Neals totalThe Neal Company wants to estimate next years return on equity (ROE) under different financial leverage ratios. Neals total
Answer #1

ROE (R) = Net income/Equity

Expected ROE (m) = sum of probability weighted ROEs

Standard deviation (SD) = {Sum of [Probability*(R-m)^2]}^0.5

Coefficient of variation (CV) = SD/m

Case 1 - Debt/Capital = 0. In this case, interest expense will be zero and Equity will be equal to Capital which is 16 million.

Formula Probability (P) EBIT | Interest (0) 0.2 0.5 EBIT- EBT 4.8 2.7 0.4 40%*EBT EBT-Tax Tax @ 40% Net income (NI) 1.92 2.88

ROE = 9.11%; Standard deviation = 5.80%; CV = 0.64

Case 2 - Debt/Capital = 10%, so Debt amount = 10%*16 = 1.6 million

Equity amount = 16 -1.6 = 14.4 million

Interest expense = 9%*1.6 = 0.144 million

| Formula Probability (P) 0.2 0.5 0.3 EBIT 4.8 2.7 0.4 9%*Debt Interest (0)| 0.144 0.144 0.144 EBIT- EBT 4.656 2.556 0.256 40

ROE = 9.53%; Standard deviation = 6.46%; CV = 0.68

Case 3 - Debt/Capital = 50%, so Debt amount = 50%*16 = 8 million

Equity amount = 16 -8 = 8 million

Interest expense = 11%*8 = 0.88 million

Formula Probability (P) EBITI 4.8 02 11%*Debt Interest (0) 0.88 0.88 0.88 EBIT-1 | 40%*EBT EBT-Tax EBT Tax @ 40% Net income (

ROE = 11.63%; Standard deviation = 11.88%; CV = 1.02

Case 4 - Debt/Capital = 60%, so Debt amount = 60%*16 = 9.6 million

Equity amount = 16 - 9.6 = 6.4 million

Interest expense = 14%*9.6 = 1.344 million

Formula Probability (p) | 02 EBIT 4.8 2.7 0.4 14%*Debt Interest 1.344 1.344 1.344 EBIT-1 | 40%*EBT EBT-Tax EBT Tax @ 40% Net

ROE = 10.18%; Standard deviation = 14.55%; CV = 1.43